Solve the system of equations if possible.
2x-3y-z=15
2x+3y+z=1
2x+3y-z=3

Answer by Alan3354(69443)

2x-3y-z=15
2x+3y+z=1
2x+3y-z=3
-------------
Use elimination
2x-3y-z=15
2x+3y+z=1
---------- Add
4x = 15
x = 4
------
2x-3y-z=15
2x+3y-z=3
----------- Add
4x - 2z = 18
16 - 2z = 18
z = -1
--------
2x+3y+z=1
8 + 3y - 1 = 1
3y = -6
y = -2
-------
Most are not this easy.